Suppose K1 and K2 have the following distribution:
Scenario Probability return K1 return K2
w(1)    0.3 -10% 10%
w(2)    0.4 0% 20%
w(3)    0.3 20% -10%


(a) Find the risk of the portfolio with w1 = 30% and w2 = 70%.
(b) Find the risk of the portfolio with w1 = 50% and w2 = 50%.
(c) Which of the portfolios above (in part (a) and (b)), has higher expected returns?
Answer #1
Probability(P) Return(%) P*Return Deviation form expected return (D1) PD^2
0.3 -10 -3 -13 50.70
0.4 0 0 -3 3.60
0.3 20 6 17 86.70

Expected Return = \sumP*Return

= -3+0+6

= 3%

Variance = \sumPD^2

= 50.7+3.6+86.7

= 141

Standard Deviation = \sqrtVariance

= \sqrt141

= 11.87%

Probability(P) Return(%) P*Return Deviation form expected return (D2) PD^2
0.3 10 3 2 1.20
0.4 20 8 12 57.60
0.3 -10 -3 -18 97.20

Expected Return = \sumP*Return

= 3+8-3

=8%

Variance = \sumPD^2

= 1.2+57.6+97.2

= 156

Standard Deviation = \sqrtVariance

= \sqrt156

= 12.49%

Probability(P) Deviation (D1) Deviation (D2) P*D1*D2
0.3 -13 2                         (7.8)
0.4 -3 12                       (14.4)
0.3 17 -18                       (91.8)

Covariance = \sumP*D1*D2

= -7.8-14.4-91.8

= -114

(a) Find the risk of the portfolio with w1 = 30% and w2 = 70%.

Standard Deviation of Portfolio = \sqrt(W1*SD1)^2+(W2*SD2)^2+(2*W1*W2*COV12)

= \sqrt(.3*11.87)^2+(.7*12.49)^2+(2*.3*.7*-114)

= \sqrt41.24077

= 6.42%

The return of a portfolio is the weighted average return of the securities which constitute the porfolio

Portfolio Return = .3*3+.7*8

= 6.50%


(b) Find the risk of the portfolio with w1 = 50% and w2 = 50%.

Standard Deviation of Portfolio = \sqrt(W1*SD1)^2+(W2*SD2)^2+(2*W1*W2*COV12)

= \sqrt(.5*11.87)^2+(.5*12.49)^2+(2*.5*.5*-114)

= \sqrt17.22425

= 4.15%

The return of a portfolio is the weighted average return of the securities which constitute the porfolio

Portfolio Return = .5*3+.5*8

= 5.50%


(c) Which of the portfolios above (in part (a) and (b)), has higher expected returns?

Portfolio in part a has higher expected return.